# Criteria for optimal entanglement-assisted long baseline imaging protocols

### v1

@@ Abstract
Quantum-entanglement-assisted telescopy protocol have been proposed and studied as a means to extend the baseline of optical interferometric telescopes. However, determining the optimality of specific quantum-assisted protocols remains an open question. We introduce a new framework for systematically characterizing different entanglement-assisted telescopy protocols by integrating tools from quantum metrology with the superselection rule (SSR) established in quantum information theory. We demonstrate that the estimation of interference fringes in a quantum telescopy can be objectively quantified and compared using the quantum Fisher information under the SSR constraint. Our framework demonstrates that interference fringe estimation in quantum telescopy can be objectively evaluated and compared using the quantum Fisher information under the SSR constraint. This approach defines the fundamental limits of astronomical estimation problems with finite entanglement resources. Building on this foundation, we propose new entanglement-assisted protocols that outperform previously known methods. Moreover, we show that these protocols are compatible with linear optical technology and could inspire the development of improved quantum telescopy schemes for near-term, lossy, repeater-less quantum networks.
@@ I Introduction
The locality constraint, as introduced in [11], is a well-recognized limitation in classical astronomical imaging operating without direct detection. Non-local resources from entangled-assisted state, however, enable overcoming this constraint. Nevertheless, the locality constraint alone is insufficient for comparing different quantum telescopy protocols. A previously overlooked yet critical constraint arises from the absence of a phase reference—formally described as the superselection rule (SSR)[19]. By examining the quantum Fisher information of astronomical photon states under the SSR constraint, we demonstrate that an SSR associated with $U(1)$ symmetry and photon number conservation[20, 21] imposes additional restrictions on the amount of information extractable from astronomical interference measurements. To address this limitation, we systematically analyze and utilize various ancilla states as a phase reference (PR) in order to effectively overcome the SSR constraint. Importantly, this framework allows us to explore the fundamental limits of ancilla-state-assisted schemes, quantify the effectiveness of different ancilla states, and propose improved entanglement-assisted protocols. Our study provides a deeper understanding of the limitations and potential of quantum telescopy protocols in general [15, 17]. Furthermore, we discuss the practical implementation of these protocols using linear optics and classical communication, offering new insights into the realization of quantum telescopy protocols in near-term, repeater-less quantum networks. This approach builds on established techniques from linear optical quantum computation (LOQC) [22]
@@ II.1 Quantum estimation problem
$\mathbb{F}\leq\mathbb{H}\leq 1/\text{Cov}$ (3)
@@
The above QFI is achievable in the ideal case, including direct interference with noise-less and loss-less optical channel, continuous variable teleportation[17] and the control-gate-based protocol [14] with perfect phase reference and quantum memory.
The optimal performance of quantum telescopy protocol discussed above is achievable either through an ideal direct interference scheme or by assuming the existence of a perfect phase reference, without addressing how such a reference can be realized. In the following, we formally define the estimation problem for quantum telescopy under the superselection rule (SSR) constraint and outline how to overcome this limitation using a phase reference in the most resource-efficient manner, which also allows us to quantify and compare different ancilla-assisted quantum telescopy protocol.
@@ II.2 Superselection rule
under the SSR constraint and effectively, a restriction will be imposed to the possible bipartite state in the system:
@@
where $P^{A(B)}_{n}=|n\rangle\langle n|$ represents projector on $H^{A}_{n}$.
@@ III Quantum Fisher Information with superselection rule
Therefore, the achievable QFI scales at most $\epsilon^{2}$, which could already explain why an intensity interferometer is not efficient in very-long baseline interferometry[30, 5], and extra resources is required if direct detection is not possible due to noise or loss in optical channel.
The extra resources required to overcome superselection rule has been known as the phase reference (PR) and has been extensively discussed in quantum information science [19]. Next, we will turn to analyze these particular phase reference resources that are needed to supply between telescopes A and B to “lift” the $U(1)-$ superselection rule, which could be in general be described by an additional ancilla state $\rho_{a}$ that acts as a phase reference for the source state $\rho_{s}$[19]. Analog to the nonlocal resources (NL) source used to overcome locality constrains, in the following, we will study the phase reference (PR) source to fight against the superselection rule constrains. Formally speaking, when an ancilla state is included, the composite system takes the following form under SSR constraint:
@@
Moreover, in order to find the optimal ancilla state that can saturate or asymptotically saturate the above inequality, it is convenient to define the following QFI ratio to quantify the performance of different ancilla states.
@@
[…]
In this work, we describe the most general ancilla state using the following notation:
@@
where we use the notation $|n_{A},m_{B}\rangle$ to represent a state with $n$ photons at $A$ and $m$ photons at $B$, without specifying the detailed mode configurations. This abstraction encompasses states for various cases, such as:
@@
The abstraction provided by this notation is particularly useful for discussing the quantum Fisher information ratio, as the mode configuration does influence the calculation, as detailed in later sections and Appendix C.
@@
Here we neglect the higher term $O(\epsilon)$ for simplicity. So $\mathbb{h}=0$ should actually be understood as $\mathbb{h}=O(\epsilon)$.
To make sense of our discussion, some constraints on the shared ancilla state $|\psi\rangle_{a}$ need to be imposed, a natural one is to limit the maximum photon number to some fixed value $N$, with this we have
@@
We leave the derivation of these upper bounds in the appendix C but would like to point out that the optimal ancilla states that saturate the above bound have the form,
@@
This indicates that the optimal ancilla state always has coefficients $f_{n,m}=\delta_{n,N-m}f_{n}$, where $N$ is the amount of resources used in the ancilla system. And the Quantum Fisher information ratio gets simplified to:
@@
CV-teleportation[17, 18] $\sum\limits_{n=0}^{\infty}\frac{\sqrt{2\langle N\rangle^{n}}}{\sqrt{(2+\langle N\rangle)^{n+1}}}|n\rangle_{A}|n\rangle_{B}$ $\geq 1-\frac{1}{\epsilon\langle N\rangle}^{\dagger}$  ✗
@@
Table 1: Summary on the QFI ratio $\mathbb{h}$ computed for different ancilla states used in the quantum telescopy protocol. The realizations of the QFI-ratios of these schemes are categorized by whether they could be achieved with local operation with classical communication (LOCC), and linear optics. The ancilla state is considered as a non-local (NL) resource, phase-reference (PR) resource (to lift the SSR), or both. Importantly, not all non-local ancilla states can be used to build a phase reference and overcome the SSR, and there exist classically correlated states, that can be used to build a phase reference, but can not help to overcome locality constraints; Therefore, a good ancilla state should be both NL resource and PR resource. : assuming an extra perfect share-phase reference exists, i.e., $|\alpha\rangle_{A}|\alpha\rangle_{B}$ with $|\alpha|\rightarrow\infty$; : with preprocessing.
We now provide explicit examples of different ancilla-assisted quantum telescopy protocols and their corresponding quantum Fisher information under SSR constraints. It is worth noting that the direct interference scheme does not belong to this category, as it is not subject to SSR constraints. The performance of different ancilla states is summarized in Table 1, and the comparison to the numerical optimal ancilla state is given in Fig. 2.
@@ IV.1 GJC scheme
Importantly, the quantum Fisher information (ratio) above is known to be achievable using correlated inference measurements (explicit protocols exist that achieve a Fisher information ratio saturating this bound). However, as noted in the original paper [10] and many subsequent discussions, such a scheme will succeed only $50\%$ of the time thereby providing a Fisher information ratio of $\frac{1}{2}$. Here we show that this $\frac{1}{2}$ ratio is not only any bound from a specific measurement protocol (the correlated interference measurement in [10] ), but also fundamental on the level of quantum fisher information when superselection rule is properly reconginized.
@@ IV.2 N-copy SPE
The discussion on N-copy-SPE state has been attempted in several works [16, 15], however, its performance has not been carefully studied yet  [^1] It was argued in [15], that the upper bound FI scales $1-\frac{1}{N+1}$, however, the conclusion is drawn assuming all photons enter the same telescopes, which could only give an loose bound scales $1-\frac{1}{2^{N}}$ Following Proposition 1, we can immediately show that the QFI-ratio using N-copy SPE is:
@@
where, we use $|n_{A},(N-n)_{B}\rangle$ to represent a state with $n$ photon at telescope $A$ and $m$ photon at telescope B for simplicity without specifying its mode configurations (i.e., it is indeed a superposition state of different mode configurations). With this, we can compute the QFI-ratio as:
@@
Agreeing with the Fisher information computed and conjectured in [16, 15]. However, this QFI cannot be achieved with all known protocols using local operations and classical communication (LOCC) and linear optics in general, which contradicts the special case discussed in [15], where only point-like sources were considered.
Similarly, in [16], another (LOCC) protocol utilizing N-copy SPE was proposed, and a Fisher information ratio $\mathsf{F}=\frac{N}{N+1}$ was computed. However, upon closer inspection of the protocol we note that it is not implementable with linear optics and requires non-linear operations. We refer the reader to the appendix 2 for a detailed discussion on this protocol.
@@ IV.3 KLM-like scheme
In the following, we introduce several novel quantum telescopy schemes inspired by the Knill-Laflamme-Milburn scheme studied in linear optical quantum computation [22]. A detailed discussion of these protocols is provided in the appendix 3.
@@
$|\psi\rangle^{\text{Sub}}_{a}=\sum\limits_{n=0}^{N}\frac{\sqrt{2}\sin(\frac{n+1}{N+2}\pi)}{\sqrt{N+2}}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
@@ IV.4 CV teleportation
At first glance, from Proposition 1, one can show that $\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]=0$, which contradicts the protocol introduced in [17, 18] but aligns with the argument made in [34]. As we will explain below, this apparent discrepancy arises from a common misunderstanding in discussions of CV quantum teleportation [35, 36], necessitating a closer examination of the actual resources employed in these protocols.
@@ IV.6 Correlated coherent state
This observation also highlights the fact that the correlated coherent state is not a nonlocal (NL) resource. Furthermore, it serves as an example illustrating that a “local scheme” cannot saturate the Quantum Fisher Information (QFI) in a nonlocal scenario, as demonstrated in other discussions [38]. Similar to the discussion in [11] and as we explain in detail in Appendix 7, we show that, in general, if $\rho_{A}^{T_{A}}>0$ (i.e., the state has a positive partial transpose), the Fisher Information (FI) ratio will always approach zero. This indicates that, in this particular example, it is the lack of nonlocality (entanglement) in the state—rather than its phase reference property—that limits its performance in the quantum telescopy.
@@
We give a short description of a conceivable linear optical implementation of our protocol to illustrate our methodology, with a more detailed derivation in the appendix E. This protocol is an idealisation, as it assumes the loss-free distribution of the ancilla state, such as provided by a repeater-network. The implementation is illustrated in Fig. 3, and operates as follows:
@@
Intuitively, this approach teleports one part of the bipartite source state from telescope A to telescope B. Since neither locality nor superselection rule is a restriction in this protocol, the optimal quantum telescope can be achieved asymptotically with linear optics. Note that the LOCC-based protocol may require a free space delay line at telescope B to compensate for the time required for classical information transmission. However, such a delay line could be made more resilient against noise in a lab compared to the optical channel interconnecting the two telescopes. It could be based, for example, on a broadband all-optical delay-line quantum memory[41] or vacuum delay line [42].
With the advent of quantum repeaters, the lossless distribution of various types of entangled states across nodes in a quantum network has become a promising possibility[43, li2024generalized, Miguel2023]. Similar to the original quantum telescopy protocol[10], a quantum repeater network addresses two key challenges: phase noise, caused by path length variations in interferometers, can be mitigated through active stabilization or entanglement distillation, while photon loss in optical channels can be managed using the repeater protocols. These advancements lay the foundation for the robust implementation of entangled-state distribution in future quantum networks and enable the reliable realization of the quantum telescopy protocols proposed in this letter.
Note that, the implementation above only works for ancilla states in Eq. 26, while the $N$-copy SPE state is not of this particular form with the same mode configurations, we discussed this issue in detail with comments on two related works[14, 15] in the Appendix 2. We want to emphasize that allthough it was argued in both papers that they have proposed linear optical implementation and achieved a Fisher information that seemingly saturates the Quantum Fisher information bound derived in the previous section, we point out some overlooked facts in both papers, see Appendix 2. Our numerical analysis indicates that the scaling of Fisher information ratio is rather $\mathbb{f}\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}1-\frac{2}{N}$ with linear optics. And this happens not because of the LOCC constraints on the measurement, but because linear optics can only become universal with extra resources[44, 45].

### v2

@@ Abstract
Quantum-entanglement-assisted telescopy protocol have been proposed and studied as a means to extend the baseline of optical interferometric telescopes. However, determining the optimality of these quantum-assisted protocols remains an open question. Here we introduce a noval framework for systematically characterizing different entanglement-assisted telescopy protocols by integrating tools from quantum metrology with the superselection rule (SSR) from quantum information theory. In our approach, the estimation of interference fringes is rigorously quantified using the quantum Fisher information under the SSR constraint. This approach defines the fundamental limits of astronomical estimation problems with finite entanglement resources, and building on this foundation, we propose new entanglement-assisted protocols that outperform previously known methods. Moreover, we show that these protocols are compatible with linear optical technology and could potentially inspire the development of improved quantum telescopy schemes for near-term, lossy, repeater-less quantum networks.
@@ I Introduction
The locality constraint, as introduced in [11], is a well-recognized limitation in classical astronomical imaging operating without direct detection. Non-local resources from entangled-assisted state, however, enable overcoming this constraint. Nevertheless, the locality constraint alone is insufficient for comparing different quantum telescopy protocols. A previously overlooked yet critical constraint arises from the absence of a phase reference—formally described as the superselection rule (SSR)[19]. By examining the quantum Fisher information of astronomical photon states under the SSR constraint, we demonstrate that an SSR associated with $U(1)$ symmetry and photon number conservation[20, 21] imposes additional restrictions on the amount of information extractable from astronomical interference measurements. To address this limitation, we systematically analyze and utilize various ancilla states as a phase reference (PR) in order to effectively overcome the SSR constraint. Importantly, this framework allows us to explore the fundamental limits of entanglement-assisted schemes, quantify the effectiveness of different ancilla states, and propose improved entanglement-assisted protocols. Our study provides a deeper understanding of the limitations and potential of quantum telescopy protocols in general [15, 17]. Furthermore, we discuss the practical implementation of these protocols using linear optics and one-way classical communication assisted by local-delay line memory, offering new insights into the realization of quantum telescopy protocols in near-term, repeater-less quantum networks that goes beyond traditional passive method[10]. This approach builds on established techniques from linear optical quantum computation [22].
@@ II.1 Quantum estimation problem
$1/\text{Cov}\leq\mathbb{F}\leq\mathbb{H}$ (3)
@@
Compared to previous works [11, 16, 17, 18], which used Fisher information as the metric for comparing various quantum telescopy protocols, this letter begins by discussing the quantum Fisher information associated with the source state $\rho_{s}^{AB}$ under specific constraints, aiming to establish a fundamental upper bound. The achievability of this quantum Fisher information will be examined through specific protocols in later sections.
@@
The above QFI is achievable in the ideal direct interference scheme with noise-less and loss-less optical channel.
Without direct detection, however, a phase reference becomes essential for realizing an optimal quantum telescopy protocol. In many previous studies, a perfect phase reference was implicitly assumed—for example, in both the continuous-variable teleportation protocol [17] and the control-gate-based protocol [14]. However, the practical realization of such a phase reference has not been adequately addressed.
In the following, we formally define the estimation problem for quantum telescopy under the superselection rule (SSR) constraint and present an efficient approach to overcoming this limitation. This method enables us to systematically analyze and compare various entangled-assisted states that serve as phase reference resources. Moreover, it provides a resource-efficient strategy for establishing a phase reference, thereby allowing us to quantitatively assess and compare all different entanglement-assisted quantum telescopy protocols.
@@ II.2 Superselection rule
and effectively, the possible bipartite state in the system under the SSR constraint has the form:
@@
where $P^{A(B)}_{n}=|n\rangle\langle n|$ represents projector on $H^{A(B)}_{n}$.
@@ III Quantum Fisher Information with superselection rule
Since the first two term contain no information about $g$, the achievable QFI scales at most $\epsilon^{2}$. This could already explain why an intensity interferometer is not efficient in very-long baseline interferometry for sensing weak astronomical source[30, 5].
The additional resources required to overcome the superselection rule are usually refered as the phase reference (PR) and have been extensively discussed in the field of quantum information science [19]. Next, we will analyze the specific phase reference resources required between telescopes A and B to ‘lift’ the $U(1)-$ superselection rule. Generally, this can be described by an additional ancilla state $\rho_{a}$, which serves as a reference for the source state $\rho_{s}$ [19]. Analogous to the nonlocal resources (NL) used to overcome locality constraints, we will subsequently explore the phase reference (PR) sources required to counter superselection rule constraints. Formally, when an ancilla state is incorporated, the composite system under the SSR constraint is structured as follows:
@@
In order to find the optimal ancilla state that can saturate or asymptotically saturate the above inequality, it is convenient to define the following QFI ratio to quantify the performance of different ancilla states.
@@
which takes value in $[0,1]$ from lemma 14. And ${\rho}^{AB}$ is given in Eq. 8. In the following, we will refer to this quantity as QFI ratio.
It is important to reiterate that the QFI-ratio serves only as an upper bound for the achievable FI-ratio in actual protocols. However, in this letter, we will discuss not only the upper bound of the QFI-ratio but also its achievability.
To get an upper bound for QFI-ratio for arbitary mode entangled state, in this work, we describe the most general ancilla state using the following notation:
@@
where we use the notation $|n_{A},m_{B}\rangle$ to represent a state with $n$ photons at $A$ and $m$ photons at $B$, without specifying the detailed mode configurations. This abstraction encompasses states with various mode entangled state, such as:
@@
The abstraction provided by this notation is useful for discussing the quantum Fisher information ratio:
@@
Here we neglect the higher term $O(\epsilon)$ for simplicity. So $\mathbb{h}=0$ should actually be understood as $\mathbb{h}=O(\epsilon)\ll 1$.
To make of our discussion and comparison more quantitative, we impose constraints on the maximum photon number to some fixed value $N$, with this we have an universal upper bound on the achievable QFI-ratio:
@@
We leave the derivation of these upper bounds in the appendix C, where the optimal ancilla states that saturate the above bound have the form,
@@
And the QFI-ratio gets simplified to:
@@
CV-teleportation[17, 18] $\sum\limits_{n=0}^{\infty}\frac{\sqrt{2\langle N\rangle^{n}}}{\sqrt{(2+\langle N\rangle)^{n+1}}}|n\rangle_{A}|n\rangle_{B}|\alpha\rangle_{A}|\alpha\rangle_{B}$ $\geq 1-\frac{1}{\epsilon\langle N\rangle}^{\dagger}$  ✗
@@
Table 1: Summary on the QFI ratio $\mathbb{h}$ computed for different ancilla states used in the quantum telescopy protocol. The realizations of the QFI-ratios of these schemes are categorized by whether they could be achieved with local operation with classical communication (LOCC), and linear optics. The ancilla state is considered as a non-local (NL) resource, phase-reference (PR) resource (to lift the SSR), or both. Importantly, not all non-local ancilla states can be used to build a phase reference and overcome the SSR, and there exist classically correlated states, that can be used to build a phase reference, but can not help to overcome locality constraints; Therefore, a good ancilla state should be both NL resource and PR resource. : with preprocessing.
In the next section, we will provide explicit examples of different ancilla-assisted quantum telescopy protocols and their corresponding quantum Fisher information under SSR constraints. It is worth noting that the direct interference scheme does not belong to this category, as it is not subject to SSR constraints. The performance of different ancilla states is summarized in Table 1, and the comparison to the numerical optimal ancilla state is given in Fig. 2.
@@ IV.1 GJC scheme
Importantly, the quantum Fisher information (ratio) above is known to be achievable using correlated inference measurements (explicit protocols exist that achieve a Fisher information ratio saturating this bound). However, as noted in the original paper [10] and many subsequent discussions, such a scheme will succeed only $50\%$ of the time thereby providing a Fisher information ratio of $\frac{1}{2}$. Here we show that this $\frac{1}{2}$ ratio is not only any bound from a specific measurement protocol (the correlated interference measurement in [10] ), but also fundamental on the level of quantum fisher information when superselection rule is imposed.
@@ IV.2 N-copy SPE
The discussion on N-copy-SPE state has been attempted in several works [16, 15], however, its performance has not been carefully studied yet. Following Proposition 1, we can immediately show that the QFI-ratio using N-copy SPE is:
@@
where, we use $|n_{A},(N-n)_{B}\rangle$ to represent a state with $n$ photon at telescope $A$ and $m$ photon at telescope B for simplicity, whichis indeed a superposition state of different mode configurations. With this, we can compute the QFI-ratio as:
@@
Agreeing with the Fisher information computed and conjectured in [16, 15]. However, this QFI cannot be achieved with all known protocols using linear optics in general, which contradicts the special case discussed in [15], where only point-like sources were considered.
Similarly, in [16], another (LOCC) protocol utilizing N-copy SPE was proposed, and a Fisher information ratio $\mathsf{F}=\frac{N}{N+1}$ was computed. However, upon closer inspection of the protocol we note that it is not implementable with linear optics and requires non-linear operations. We refer the reader to the appendix D.2 for a detailed discussion on this protocol.
@@ IV.3 KLM-like scheme
In the following, we introduce several novel quantum telescopy schemes inspired by the Knill-Laflamme-Milburn scheme studied in linear optical quantum computation [22]. A detailed discussion of these protocols is provided in the appendix D.3.
@@
$|\psi\rangle^{\text{Opt}}_{a}=\sum\limits_{n=0}^{N}\frac{\sqrt{2}\sin(\frac{n+1}{N+2}\pi)}{\sqrt{N+2}}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
@@ IV.4 Continuous-variable protocol
At first glance, from Proposition 1, one can show that $\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]=0$, which contradicts the protocol introduced in [17, 18]. As we will explain below, this apparent discrepancy arises from a common misunderstanding in the discussions of CV quantum teleportation [33, 34], which requires a closer examination of the actual resources employed in these protocols.
@@ IV.6 Correlated coherent state
This observation also highlights the fact that the correlated coherent state is not a nonlocal (NL) resource. Furthermore, it serves as an example illustrating that a “local scheme” cannot saturate the Quantum Fisher Information (QFI) in a nonlocal scenario, as demonstrated in other discussions [36]. Similar to the discussion in [11] and as we explain in detail in Appendix D.7, we show that, in general, if $\rho_{A}^{T_{A}}>0$ (i.e., the state has a positive partial transpose), the Fisher Information (FI) ratio will always approach zero. This indicates that, in this particular example, it is the lack of nonlocality (entanglement) in the state—rather than its phase reference property—that limits its performance in the quantum telescopy.
@@
We give a short description of a conceivable linear optical implementation of our protocol to illustrate our methodology, with a more detailed derivation in the appendix E. This protocol is an idealisation, as it assumes the loss-free distribution of the ancilla state, such as provided by a repeater-network, moreover, it differs from the traditional passive scheme, where classical communication is completely forbidden[10, 15]. The implementation is illustrated in Fig. 3, and operates as follows:
@@
Intuitively, this approach teleports one part of the bipartite source state from telescope A to telescope B. Since neither locality nor superselection rule is a restriction in this protocol, the optimal quantum telescope can be achieved asymptotically with linear optics. Note that the one-way LOCC-based protocol may require a free space delay line at telescope B to compensate for the time required for classical information transmission. However, such a delay line could be made more resilient against noise in a lab compared to the optical channel interconnecting the two telescopes. It could be based, for example, on a broadband all-optical delay-line quantum memory[39] or local vacuum beam guide [40].
With the advent of quantum repeaters, the lossless distribution of various types of entangled states across nodes in a quantum network has become a promising possibility[41, li2024generalized, Miguel2023]. Similar to the original quantum telescopy protocol[10], a quantum repeater network addresses two key challenges: phase noise, caused by path length variations in interferometers, can be mitigated through active stabilization or entanglement distillation, while photon loss in optical channels can be managed using the repeater protocols. These advancements lay the foundation for the robust implementation of entangled-state distribution in future quantum networks and enable the reliable realization of the quantum telescopy protocols proposed in this letter.
Note that, the implementation above only works for ancilla states in Eq. 26, while the $N$-copy SPE state is not of this particular form with the same mode configurations, we discussed this issue in detail with comments on two related works[14, 15] in the Appendix D.2. We want to emphasize that allthough it was argued in both papers that they have proposed linear optical implementation and achieved a Fisher information that seemingly saturates the Quantum Fisher information bound derived in the previous section, we point out some overlooked facts in both papers, see Appendix D.2.